x = COST TT 7) Find the slope of the line tangent to at t = y = 8 sint 2 1 - 3 8) Find the length of 0<t<1 1 - 21
Answer #1

> x=20$CH) ,y=8 Sinct) data-Sinct), y = 860&C+ at t= 5/2 dx=-1 de = 0 ..slope af tangent line be mzdy - dylat dr = / drrat m=